(a) If the molar solubility of Ca3(PO4)2 at 25 oC is 4.53e-08 mol/L, what is the Ksp at this temperature?

Ksp =




(b) It is found that 8.70e-09 g of Ga(OH)3 dissolves per 100 mL of aqueous solution at 25 oC. Calculate the solubility-product constant for Ga(OH)3.

Ksp =




(c) The Ksp of Ag2C2O4 at 25 oC is 5.40e-12. What is the molar solubility of Ag2C2O4?

solubility =  mol/L
Answer #1

a)

At equilibrium:

Ca3(PO4)2 <----> 3 Ca2+ + 2 PO43-

   3s 2s

Ksp = [Ca2+]^3[PO43-]^2

Ksp = (3s)^3*(2s)^2

Ksp = 108(s)^5

Ksp = 108(4.53*10^-8)^5

Ksp = 2.06*10^-35

Answer: 2.06*10^-35

b)

s = 8.70*10^-9 g / 100 mL

= 8.70*10^-9 g / 0.100 L

= 8.70*10^-8 g / L

Molar mass of Ga(OH)3,

MM = 1*MM(Ga) + 3*MM(O) + 3*MM(H)

= 1*69.72 + 3*16.0 + 3*1.008

= 120.744 g/mol

Molar mass of Ga(OH)3= 120.744 g/mol

s = 8.7*10^-8 g/L

To covert it to mol/L, divide it by molar mass

s = 8.7*10^-8 g/L / 120.744 g/mol

s = 7.205*10^-10 mol/L

At equilibrium:

Ga(OH)3 <----> Ga3+ + 3 OH-

   s 3s

Ksp = [Ga3+][OH-]^3

Ksp = (s)*(3s)^3

Ksp = 27(s)^4

Ksp = 27(7.205*10^-10)^4

Ksp = 7.277*10^-36

Answer: 7.28*10^-36

c)

At equilibrium:

Ag2C2O4 <----> 2 Ag+ + C2O42-

   2s s

Ksp = [Ag+]^2[C2O42-]

5.4*10^-12=(2s)^2*(s)

5.4*10^-12= 4(s)^3

s = 1.105*10^-4 M

Answer: 1.10*10^-4 M